# OmniPred: Language Models as Universal Regressors

### v3

@@ Abstract
Over the broad landscape of experimental design, regression has been a powerful tool to accurately predict the outcome metrics of a system or model given a set of parameters, but has been traditionally restricted to methods which are only applicable to a specific task. In this paper, we propose OmniPred, a framework for training language models as universal end-to-end regressors over $(x,y)$ evaluation data from diverse real world experiments. Using data sourced from Google Vizier, one of the largest blackbox optimization databases in the world, our extensive experiments demonstrate that through only textual representations of mathematical parameters and values, language models are capable of very precise numerical regression, and if given the opportunity to train over multiple tasks, can significantly outperform traditional regression models.
[…]
@@ 1 Introduction
Figure 1: Overview of our method. Using heterogenous offline blackbox function evaluations collected from systems optimized using Google Vizier, we train a LM-based regressor.
In recent years, large language models (LLMs) have emerged as powerful tools for processing textual representations at scale over massive heterogeneous datasets to represent complex relationships between input features and output labels. Given that LLMs have been shown to be effective for a variety of tasks beyond natural language processing, such as coding (Li et al., 2022), symbolic mathematics (Lewkowycz et al., 2022), and scientific reasoning (Singhal et al., 2022), it is reasonable to wonder: Can language models be used for regression?
@@
  • To the best of our knowledge, we propose OmniPred, the first scalable yet simple metric prediction framework based on constraint-independent textual representations, applicable to general input spaces.
@@
  • By simultaneously multi-task learning across vastly different input spaces and objectives, in many cases OmniPred can outperform traditional regression models such as MLPs and boosted trees.
  • These transfer learning benefits persist even on unseen tasks after locally finetuning OmniPred on small amounts of new evaluation data.
@@ 2 Related Work and Motivation
Traditional regression methods have widely used statistical techniques such as Gaussian Processes (GPs), tree-based methods, and multilayer perceptrons (MLPs), to predict a scalar objective given a fixed-length feature vector, commonly seen in tabular data settings. Multitask (Bonilla et al., 2007) and contextual (Krause & Ong, 2011) variants have been further proposed for transfer learning purposes, but still require fixed-length tensor representations of $x$, and can thus only use previous $x$ from the same input space. Additional recent works utilizing deep learning-based regressors include Transformers (Hollmann et al., 2023; Huang et al., 2020), recurrent neural networks (Hashemi et al., 2018), graph neural networks (Lukasik et al., 2020; Gao et al., 2023), and deep-hierarchical GPs (Fan et al., 2024), which allow length-independence. Even so, a frequent issue is still the reliance on tensor representations of $(x,y)$.
Tensor representations are commonly constraint-dependent, where each tensor element may need to be in a reasonable numerical range (e.g. in $[-1,1]$) as inputs to a model. Thus to represent $x$, every categorical feature must be one-hot embedded against user-provided choices, and scalar features may need to be rescaled against user-provided bounds. Dynamic yet minor input space changes such as new bounds or additional categories, are incompatible with this static representation. To represent $y$, a raw objective in $\mathbb{R}$ may also need to be rescaled, which can be problematic at test-time when encountering outlier $y$-values. Dealing with this issue leads to implementing complicated nonlinear warpings (Daimon, 2011; Yeo & Johnson, 2000), many of which are also data-dependent (e.g. require storing min/max values from training data).
@@
In principle, an ideal regressor should process $x$ and output $y$, both in absolute terms, independent of changing external statistics or search constraints. For example, if the objective is $f(x)=\exp(x)$, then the regressor’s prediction for $f(2)$ should be invariant regardless if the constraint is $x\in[1,5]$ or $x\in[0,100]$. One way to accomplish this is via token based representations of the data, which is instead parsed by tokens or symbols discretely (Zhou et al., 2023). This immediately unlocks a large amount of transferrability when dealing with variable-length inputs and additional contextual metadata.
This token-based paradigm has shown great success in the case of reinforcement learning from human feedback (Ziegler et al., 2019), where regression over textual responses (the “$x$”), also known as reward modelling, has been crucial to the success of recent interactive LLMs such as ChatGPT (OpenAI, 2022) and Bard (Thoppilan et al., 2022). Here, LLMs are able to imitate human ratings in the form of pairwise rankings (the “$y$”) or probabilistic scores $y\in[0,1]$ (Bradley & Terry, 1952).
While the overwhelming current focus has been on subjective human-based feedback needed for determining aspects such as creativity, safety, and personality, much less attention has been given towards language models for evaluating complex and natural systems common to experimental design, which consist of much more objective and numeric-based data where $y\in\mathbb{R}$. Given multiple works which have shown the brittleness and unreliability of numerical processing in language models (Hendrycks et al., 2021; Nogueira et al., 2021), it is not immediately obvious that language models are capable of high-precision numerical prediction over token-based representations. This is a crucial technical challenge which our paper resolves in the quest for a general-purpose predictor.
@@ 3.1 Preliminaries and Problem Definition
Based on standard blackbox optimization terminology (Golovin et al., 2017; Liaw et al., 2018), for a given task $\mathcal{T}=(\mathcal{X},f,\mathcal{D},m)$, we assume there is an inherent objective function $f:\mathcal{X}\rightarrow\mathbb{R}$ for which we obtain trials $(x,y)$ from evaluating suggestions $x$, selected from a (possibly implicit) input space $\mathcal{X}$. We define a study as an offline collection of trials $\mathcal{D}=\{(x_{1},y_{1}),...,(x_{T},y_{T})\}$. To distinguish between different tasks, there may be observable task-level metadata $m$, which can additionally characterize the task and potentially even describes the behavior of the corresponding objective $f(x)$.
The goal of standard metric prediction is to obtain a distribution function $s:\mathcal{X}\rightarrow\mathcal{P}(\mathbb{R})$ such that $s(x)$ accurately approximates $f(x)$ over a distribution of inputs from $\mathcal{X}$, provided that a training set $\mathcal{D}^{train}$ is given. In our particular case, we also provide our language model with multi-task training data $\cup\{\mathcal{D}_{1}^{train},\mathcal{D}_{2}^{train},...\}$ from other tasks $\{\mathcal{T}_{1},\mathcal{T}_{2},...\}$. While these extraneous tasks do not contain exact evaluations over the $f$ at hand and may even have different input spaces, training on such additional extraneous data may still lead to transferrability, especially for similar tasks.
A common and unified way to measure the accuracy of predictors (deterministic or stochastic) is to compute the gap between a final pointwise prediction against the true objective value, using a regressor-dependent aggregator $\alpha:\mathcal{P}(\mathbb{R})\rightarrow\mathbb{R}$ such as median or mean. However, since different studies can have vastly different objective scales (e.g. CIFAR10 accuracies are within $[0,1]$ while synthetic objectives are within $[10^{2},10^{9}]$), we must therefore normalize the difference based on per-study statistics, i.e. for a specific task, we define the study error as a normalized mean absolute error (MAE):
$\frac{1}{y_{\max}-y_{\min}}\frac{1}{|\mathcal{D}^{test}|}\sum_{(x,y)\in\mathcal{D}^{test}}\lvert\alpha(s(x))-y\rvert$ (1)
@@ 3.2 Language Model
In this paper, we focus on the standard multi-task regression setting, in which for a given trial $(x,y)$ and task-level metadata $m$, the prompt is $(x,m)$ and response is $y$, compatible with both encoder-decoder and decoder-only settings. For simplicity, we train a standard 200M parameter T5 encoder-decoder (Raffel et al., 2020) from scratch. We wish to learn a single set of weights $\theta$, which can be used to form a predictor $s_{\theta}(\cdot)$ given any arbitrary task $\mathcal{T}$. In contrast to settings such as (1) traditional regression requiring training a separate model $\theta_{t}$ for each task $\mathcal{D}_{t}$ or (2) requiring completely evaluated trajectories over specialized trial tokenizations for in-context learning (Chen et al., 2022; Hollmann et al., 2023), our setting maximizes the usage of training data, much of which may contain unfinished trajectories or non-standard $x$-formats.
Representation: To facilitate training a single model over multiple heterogeneous studies, a crucial guiding principle as mentioned before, is to express $x$ and $y$ in absolute fashion, independent of the input space and numeric scaling of the specific study. Thus, we express $x$ in a key-value format, directly mapping parameter names to values, but do not represent [^1] In applications such as code search, it is even infeasible to express the space of all possible programs. the input space $\mathcal{X}$, allowing generalizability to conditional parameters and dynamic constraints. We represent $y$ with fixed-length custom tokens to guarantee proper decoding via token logit restriction, using specific tokens to express sign, exponent, and significant digits. Ablations over different tokenizations are conducted in Appendix A.1.
@@
$x$ batch_size:128,kernel:’rbf’,learning_rate:0.5,model:’svm’,optimizer:’sgd’
$m$ title:’classification’,user:’some-person’,description:’spam detection’, objective:’accuracy’
$y$ <+><1><2><3><E-2>
Table 2: Textual representations used for OmniPred. <normal-$\ast$> represents a single custom token. Input space and $x$ is the same as in Figure 2. Example $y$ tokenization represents a value of 1.23.
Training: To maintain simplicity and normalization-independence, we minimize standard cross-entropy loss over the $y$-value tokens, consistent with typical language model training. The model will thus implicitly learn numeric distances from training data.
Sampling and Decoding: Through regular temperature decoding, we can repeatedly sample $\hat{y}\sim s_{\theta}(x)$, to approximate the underlying distribution defined over $\mathbb{R}$. To remain robust to strong outliers, we aggregate samples using the empirical median, with ablations over different aggregation methods in Appendix A.2. Since the model may need to predict over unseen regions of the input space, we can also assess the model’s uncertainty by observing the concentration of sampled $\hat{y}$ and additionally specific log probabilities across every decoded token.
Local Finetuning: To adapt to an unseen task $\mathcal{T}_{u}$, the model can further be quickly finetuned over the tasks’s corresponding training data $\mathcal{D}^{train}_{u}$. This is a common requirement during regressor-guided search, where online function evaluations should be taken into account, e.g. in Wistuba & Grabocka (2021). Finetuning may also help to refocus over seen data, when the model is not fully optimized against a specific study, e.g. if the pretraining dataset was too large.
@@ 4 Data
4.1 Vizier Format
The abstractions in Section 3.1 above are concretely implemented in Open Source Vizier (Song et al., 2022), a research interface for blackbox and hyperparameter optimization. Every space $\mathcal{X}$ is defined by a list of parameters, each of type DOUBLE, INTEGER, DISCRETE, or CATEGORICAL with an associated value set. Every parameter may also potentially be a child parameter, only active when the corresponding parent parameter is a specific value (e.g. “beta” is active only if a parent categorical parameter selects “Adam”, but not “SGD”). An example is shown in Figure 2.
Figure 2: Common example of a (possibly nested) space and suggestions $x$ in Google Vizier.
Task-level metadata $m$ consists of a title, owner username, description, objective name, and optional free-form text. Since the Vizier API is meant to provide an optimization service for users, there can be many sources of transferrability due to user-specific settings. These include:
@@ 4.2 Datasets
Table 3: Relevant statistics on the Google Vizier database. We provide order estimates as there may be numerous ways to define e.g. “legitimate” studies or trials. See Appendix D for further details.
Real World Data: To investigate metric prediction over real world data which contain a rich variety of tasks, we naturally will use the database from Google Vizier (Golovin et al., 2017). Because we are not constrained to only training over fully completed trajectories over flat input spaces, our data usage is much larger than the 750K studies for training OptFormer (Chen et al., 2022), as seen from Table 3.
Since Vizier only acts as a lightweight service for blackbox optimization, we for the most part do not have online access to an actual objective $f(x)$, rather only data samples $\mathcal{D}$, and thus we must evaluate our predictor’s accuracy via a test set $\mathcal{D}^{test}\subset\mathcal{D}$. We thus need to take into account how much $\mathcal{D}_{train}$ sufficiently covers the space $\mathcal{X}$, which affects the difficulty of achieving high accuracy on the task. Influencing factors include:
@@
  • Space size: Approximate cardinality of a space $\mathcal{X}$ is $\exp(\text{parameter count})$, and thus large input spaces will naturally be less explored.

### v4

@@ Abstract
Regression is a powerful tool to accurately predict the outcome metric of a system given a set of parameters, but has traditionally been restricted to methods which are only applicable to a specific task. In this paper, we propose OmniPred, a framework for training language models as universal end-to-end regressors over $(x,y)$ data from arbitrary formats. Using data sourced from Google Vizier, one of the largest proprietary blackbox optimization databases in the world, our extensive experiments demonstrate that language models are capable of very precise numerical regression using only textual representations of mathematical parameters and values, and if given the opportunity to train at scale over multiple tasks, can significantly outperform traditional regression models.
@@ 1 Introduction
Figure 1: Overview of our method. Using heterogenous offline $(x,y)$ evaluation data collected from a variety of sources, we train a LM-based regressor.
In recent years, large language models (LLMs) have emerged as powerful tools for processing textual representations at scale over massive heterogeneous datasets to represent complex relationships between input features and output labels. Given that LLMs have been shown to be effective for a variety of tasks beyond natural language processing, such as coding (Li et al., 2022), symbolic mathematics (Lewkowycz et al., 2022), and scientific reasoning (Singhal et al., 2022), it is reasonable to wonder: Can language models be used for numeric regression?
@@
  • To the best of our knowledge, we propose OmniPred, the first scalable yet simple metric prediction framework based on free-form textual representations, applicable to general input spaces.
@@
  • By leveraging multi-task learning across vastly different input spaces and objectives, in many cases OmniPred can outperform traditional regression models such as MLPs and boosted trees.
  • These transfer learning benefits persist even on unseen tasks after online finetuning OmniPred on small amounts of new evaluation data.
@@ 2 Related Work and Motivation
Traditional regression methods have widely used statistical techniques such as Gaussian Processes (GPs), tree-based methods, and multilayer perceptrons (MLPs), to predict a scalar objective given a fixed-length feature vector, commonly seen in tabular data settings. Multitask (Bonilla et al., 2007) and contextual (Krause & Ong, 2011) variants have been further proposed for transfer learning purposes, but still require fixed-length tensor representations of $x$, and can thus only use previous $x$ from the same input space. Additional recent works utilizing deep learning-based regressors include Transformers (Hollmann et al., 2023; Huang et al., 2020; Garg et al., 2022), recurrent neural networks (Hashemi et al., 2018), graph neural networks (Lukasik et al., 2020; Gao et al., 2023), and deep-hierarchical GPs (Fan et al., 2024), which allow length-independence. Even so, a frequent issue is still the reliance on tensor representations of $(x,y)$.
Tensor representations are commonly problem-dependent, where each tensor element may need to be in a reasonable numerical range (e.g. in $[-1,1]$) as inputs to a model. Thus to represent $x$, every categorical feature must be one-hot embedded against user-provided choices, and scalar features may need to be rescaled against user-provided bounds. Dynamic yet minor input space changes such as new bounds or additional categories, are incompatible with this static representation. To represent $y$, a raw objective in $\mathbb{R}$ may also need to be rescaled, which can be problematic at test-time when encountering outlier $y$-values. Dealing with this issue leads to implementing complicated nonlinear warpings (Daimon, 2011; Yeo & Johnson, 2000), many of which are also data-dependent (e.g. require storing min/max values from training data).
@@
These issues are summarized in Table 1. In principle, an ideal regressor should process $x$ and output $y$, both in absolute terms, independent of changing external statistics or search constraints. For example, if the underlying relationship is $y=x^{2}$, then the regressor’s prediction at $x=2$ should be invariant regardless if the constraint is $x\in[1,5]$ or $x\in[0,100]$. One way to accomplish this is to represent $x$ with an independent universal tokenizer without problem-specific numeric transformations (Zhou et al., 2023). This immediately unlocks a large amount of transferrability when dealing with variable-length inputs and additional contextual metadata.
Previous works in the token-based, or effectively text-to-text paradigm, have mostly been in reinforcement learning from human feedback (Ziegler et al., 2019), where regression over textual responses (the “$x$”), also known as reward modelling, has been crucial to the success of recent interactive LLMs such as ChatGPT (OpenAI, 2022) and Gemini (Google, 2024). While such works have shown that LLMs are able to imitate human ratings in the form of ordinal variables and their softmax probabilities (Bradley & Terry, 1952), they have not shown whether LLMs are capable of regression over precise numeric-based data where $y\in\mathbb{R}$.
This is because the overwhelming current focus has been on subjective human-based feedback needed for determining aspects such as creativity, safety, and personality, which contain high aleatoric uncertainty and do not require high-precision measurements. Much less attention has been given towards feedback from complex and natural systems common to experimental design. Given multiple works (Hendrycks et al., 2021; Nogueira et al., 2021) demonstrating their brittle and unreliable numeric abilities, it is non-obvious that language models are capable of high-precision numerical prediction over token-based representations. This is a crucial technical challenge which our paper resolves in the quest for a general-purpose predictor.
@@ 3.1 Preliminaries and Problem Definition
Since regression is commonly used in blackbox optimization settings, we adopt standard terminology (Golovin et al., 2017; Liaw et al., 2018) from the field. For a given task $\mathcal{T}=(\mathcal{X},f,\mathcal{D},m)$, we assume there is a mapping $f:\mathcal{X}\rightarrow\mathbb{R}$ for which we obtain trials $(x,y)$ from evaluating an input $x$, selected from a (possibly implicit) input space $\mathcal{X}$. We define a study as an offline collection of trials $\mathcal{D}=\{x_{s},y_{s}\}_{s=1}^{T}$. To distinguish between different tasks, there may be observable task-level metadata $m$, which can additionally characterize the task and potentially even describes the behavior of the corresponding mapping $f(\cdot)$.
The goal of standard regression is to obtain a distribution mapping $s:\mathcal{X}\rightarrow\mathcal{P}(\mathbb{R})$ such that $s(x)$ accurately approximates $f(x)$ over some distribution of inputs $x\in\mathcal{X}$, provided that a training set $\mathcal{D}^{train}$ is given. In our particular case, we also provide our language model with multi-task training data $\cup\{\mathcal{D}_{1}^{train},\mathcal{D}_{2}^{train},...\}$ from other tasks $\{\mathcal{T}_{1},\mathcal{T}_{2},...\}$. While these extraneous tasks contain different objectives $f_{1},f_{2},\ldots$ and may even have different input spaces from each other, training on such additional extraneous data may still lead to transferrability, especially for similar tasks.
A common way to measure the accuracy of predictors (deterministic or probabilistic) is to compute the gap between a final pointwise prediction against the true objective value $y$. For probabilistic regressors $s:\mathcal{X}\rightarrow\mathcal{P}(\mathbb{R})$, we may aggregate by e.g. taking the median or mean of the distribution. Since different studies can have vastly different objective scales (e.g. CIFAR10 accuracies are within $[0,1]$ while synthetic objectives are within $[10^{2},10^{9}]$), we must therefore normalize the difference based on per-study statistics, i.e. for a specific task, we define the study error as a normalized mean absolute error (MAE):
$\frac{1}{y_{\max}-y_{\min}}\frac{1}{|\mathcal{D}^{test}|}\sum_{(x,y)\in\mathcal{D}^{test}}\lvert\text{Aggregate}(s(x))-y\rvert$ (1)
@@ 3.2 Language Model
We use a standard language model in which the model observes a prompt and decodes a response. For the regression setting, naturally these correspond to $x$ and $y$ respectively. However, in order to allow multi-task training, the task-specific metadata $m$ must be appended to the prompt in order to distinguish between different tasks, and thus for a given task, the prompt is actually $(x,m)$.
For simplicity, we train a relatively small 200M parameter T5 encoder-decoder (Raffel et al., 2020) from scratch to avoid any confounding effects from typical generative language pre-training. We wish to learn a single set of weights $\theta$, which can be used to form a stochastic regressor $s_{\theta}:\mathcal{X}\rightarrow\mathcal{P}(\mathbb{R})$ given any arbitrary task $\mathcal{T}$. In contrast to settings such as (1) traditional regression requiring training a separate model $\theta_{i}$ for each task $\mathcal{T}_{i}$ or (2) requiring completely evaluated trajectories over specialized $x$-tokenizations for in-context learning (Chen et al., 2022; Hollmann et al., 2023), our setting maximizes the usage of training data, much of which may contain unfinished trajectories or free-form $x$-formats.
Representation: As mentioned, since input spaces $\mathcal{X}$ and $y$-scales may vary wildly across different tasks, multi-task training requires $x$ and $y$ to be represented in absolute formats independent of the input space and numeric scaling of the specific study. Thus, we express $x$ in a key-value format, directly mapping parameter names to values, but do not represent [^1] In applications such as code search, it is even infeasible to express the space of all possible programs. the input space $\mathcal{X}$, allowing generalizability to conditional parameters and dynamic constraints. We represent $y$ with custom tokens to guarantee proper decoding via constrained decoding, using specific tokens to express sign, exponent, and significant digits, although more sophisticated alternatives (Koo et al., 2024) are available. An example is illustrated in Table 2. Ablations over different tokenizations are conducted in Appendix A.1.
@@
$x$ batch_size:128,kernel:‘rbf’,learning_rate:0.5,model:‘svm’,optimizer:‘sgd’
$m$ title:‘classification’,user:‘some-person’,description:‘spam detection’, objective:‘accuracy (%)’
$y$ <+><7><2><5><E-1>
Table 2: Textual representations used for OmniPred. $\ast$> represents a single custom token. Input space and $x$ is the same as in Figure 2. Example $y$-tokenization represents a value of $72.5725\times 10^{-1}=$ 72.5.
Training: We apply standard Prefix-LM training (Raffel et al., 2020), in which for a given (prompt, response) pair, cross-entropy losses are only computed over response tokens. Pairs are sampled from training data over multiple tasks. One could additionally make the loss more metric-aware by weighting specific tokens or even reinforce with non-differentiable scores, although we maintain simplicity in this paper for now by using uniform cross-entropy. Thus the model will implicitly learn numeric distances from training data.
Sampling and Decoding: Through regular temperature decoding, we can repeatedly sample $\widehat{y}\sim s_{\theta}(x)$, to approximate the prediction distribution over $\mathbb{R}$. To remain robust to strong outliers, instead of using empirical mean, our $\text{Aggregate}(\cdot)$ function is defined to be the empirical median, since we found it leads to lower error from ablations in Section 6.1. Since the model may need to predict over unseen regions of the input space, we can also assess the model’s uncertainty by observing the concentration of sampled predictions $\widehat{y}$ and additionally specific log probabilities across every decoded token.
Online Finetuning: To adapt to an unseen task $\mathcal{T}_{u}$ common for Bayesian optimization settings (Wistuba & Grabocka, 2021), the model can further be quickly finetuned online over the tasks’s corresponding training data $\mathcal{D}^{train}_{u}$, optionally using LoRA (Hu et al., 2022). Finetuning may also help to refocus over seen data, when the model is not fully optimized against a specific study, e.g. if the pretraining dataset was too large.
Appendix B contains additional implementation details such as the specific architecture and vocabulary used.
@@ 4 Data
Many industries possess large collections of metric data from experiments or systems. However, such data is typically not open-sourced as official training datasets for research. A natural dataset to use may come from multiple hyperparameter optimization trajectories, which tend to have $(x,y)$ evaluations from expensive experiments, expressed as blackbox objectives $y=f(x)$.
4.1 OSS Vizier Format
The abstractions in Section 3.1 above are concretely implemented in Open Source (OSS) Vizier (Song et al., 2022), a research interface for blackbox and hyperparameter optimization. Every space $\mathcal{X}$ is a Cartesian product of parameters, each of type DOUBLE (bounded continuous interval), INTEGER (bounded integer set), DISCRETE (ordered set of real numbers), or CATEGORICAL (unordered set of strings). Every parameter may also have child parameters, only active when the corresponding parent parameter is a specific value (e.g. in Figure 2, “beta” is active only if a parent categorical parameter selects “Adam”, but not “SGD”). An $x\in\mathcal{X}$ can be expressed as a tabular feature, allowing baseline comparisons against traditional regression models.
Figure 2: Common example of a (possibly nested) space and suggestions $x$ in OSS Vizier.
Task-level metadata $m$ consists of a title, owner username, description, objective name, and optional free-form text. Since the OSS Vizier API is meant to provide an optimization service for users, there can be many sources of transferrability due to user-specific settings. These include:
@@ 4.2 Datasets
Table 3: Relevant statistics on the real world database. We provide order estimates as there may be numerous ways to define e.g. “legitimate” studies or trials. See Appendix D for further details and data breakdown.
Real World Data: To investigate metric prediction over real world data which contain a rich variety of tasks, we will use real world data from Google Vizier (Golovin et al., 2017). Because we are not limited to training on fully completed trajectories over flat input spaces, we can train on more data than just the 750K studies for training the OptFormer (Chen et al., 2022), as seen from Table 3.
Since the offline dataset is collected from users’ blackbox optimization trajectories, we for the most part do not have online access to an actual objective $f(x)$, rather only data samples $\mathcal{D}$, and thus we must evaluate our predictor’s accuracy via a test set $\mathcal{D}^{test}\subset\mathcal{D}$. Thus, we need to take into account how much $\mathcal{D}_{train}$ sufficiently covers the space $\mathcal{X}$, which affects the difficulty of achieving high accuracy on the task. Influencing factors include:
@@
  • Space size: Approximate cardinality of a space $\mathcal{X}$ is exponential with respect to parameter count, and thus large input spaces will naturally be less explored.